OpenStudy (anonymous):

Evaluate the function at given values f(x) = 3-3x When x= -1,10?

OpenStudy (anonymous):

i believe this is 2 questions f(-1) = 3 - 3( -1) = 6 and f(10) = 3 - 3 (10) = - 27